# Grammar and Gameplay-aligned RL for Game Description Generation with LLMs

### v1

@@ Abstract
Game Description Generation (GDG) is the task of generating a game description written in a Game Description Language (GDL) from natural language text. Previous studies have explored generation methods leveraging the contextual understanding capabilities of Large Language Models (LLMs); however, accurately reproducing the game features of the game descriptions remains a challenge. In this paper, we propose reinforcement learning-based fine-tuning of LLMs for GDG (RLGDG). Our training method simultaneously improves grammatical correctness and fidelity to game concepts by introducing both grammar rewards and concept rewards. Furthermore, we adopt a two-stage training strategy where Reinforcement Learning (RL) is applied following Supervised Fine-Tuning (SFT). Experimental results demonstrate that our proposed method significantly outperforms baseline methods using SFT alone.
@@ I Introduction
Recently, there has been increased interest in Game Description Generation (GDG) [4, 6, 7, 8]. GDG focuses on generating game descriptions from natural language texts, making it easier for non-experts to participate in game design. In GDG tasks, In-Context Learning (ICL) with Large Language Models (LLMs) [9] has shown great promise. LLMs excel in understanding textual contexts and can perform tasks with limited domain knowledge based on a few examples provided in prompts. For example, LLMGG [10] demonstrated that enriching prompts with explanations of GDL notation and examples of game descriptions improves the accuracy of generated outputs. Additionally, GGDG [11] proposed an iterative decoding method guided by GDL grammar rules, significantly enhancing grammatical correctness. However, the iterative improvements in GGDG have restricted grammatical accuracy and do not consider the actual gameplay behavior and features obtained through simulation.
@@ II-A Game Description Language
Game Description Language (GDL) is a specialized language for describing specific games. In 2005, GGP-GDL [1] was introduced for General Game Playing, aiming to develop artificial intelligence agents capable of adapting to various games. Video Game Description Language (VGDL) [2] has been developed to represent rules and levels of 2D sprite-based games, currently modeling as many as 195 different games. Regular Boardgames (RBG) [3] is another language that combines high-level language features with low-level descriptions, enabling the representation of complex board games. The Ludi system [4], incorporating evolutionary game design, successfully led to the development of the commercially successful game ”Yavalath.” Moreover, Ludii [5] is a system adopting the ”ludemic approach,” allowing the decomposition and description of game components at a conceptual level. This enables Ludii to represent over 1,000 traditional games, including board games, card games, dice games, and tile games. Given the broad representational capability and versatility of Ludii GDL, we primarily use Ludii GDL for our research.
@@ II-B Large Language Models in Games
Since the emergence of ChatGPT [20], large language models (LLMs) have garnered significant attention, prompting exploration into diverse applications within the field of game AI [21, 22]. For example, several methods [23, 24] have been proposed for generating 2D tile-based game levels by fine-tuning GPT-2 [9]. Additionally, research is progressing on prompt-based LLM methods for level generation [25] and quest generation in role-playing games [26]. Dreamcraft [27] further demonstrates a technique to create 3D game objects within Minecraft from textual prompts. Li et al. [28] introduces a virtual pet game that achieves real-time gameplay even with smaller LLMs by employing a domain-specific distillation approach. GAVEL [29] employs an LLM fine-tuned on Ludii game descriptions as a mutation operator in evolutionary search, aiming to create novel games. The objective of GAVEL differs from our research, which focuses on generating game descriptions that maintain consistency with natural language text. LLMGG [10] proposes a method for generating both rules and levels in VGDL using LLMs. LLMaker [30] improves content consistency by utilizing function calling, but its creativity is limited to the scope of the defined functions. GGDG [11] utilizes iterative decoding based on the grammar of Ludii GDL to enhance the grammatical correctness of generated game descriptions. In our study, we aim to enhance LLMs through reinforcement learning to bring the game characteristics of generated game descriptions closer to ground truth.
@@ II-C RL for LLMs
RL-based fine-tuning of LLMs [31, 14, 13, 15] has been attracting attention. Reinforcement Learning with Human Feedback (RLHF) [31] aligns model behaviors with human preferences by using human feedback as rewards. Recent methods [14, 13] have explored accuracy-based reward functions without human feedback, notably improving logical reasoning tasks such as mathematics and programming. RL-finetuned models are known to exhibit enhanced reasoning capabilities [14]. Additionally, advancements in RL algorithms led to the development of Group Relative Policy Optimization (GRPO) [32], a variant of Proximal Policy Optimization (PPO) [33] optimized for fine-tuning LLMs. In this study, we focus on game description generation and define a reward function based on both grammatical correctness and conceptual relevance to the game in generated descriptions, employing GRPO to fine-tune the LLM.
Figure 2: An example of game description generation for the game Tic-Tac-Toe. $x$ is text that explains games in natural language. $y$ is a game description in Ludii GDL, a Game Description Language.
@@ III Problem Setting
In this section, we define Game Description Generation (GDG), the primary task addressed in this paper. As shown in Fig. 2, the task involves providing a query $x$ to a large language model (LLM), which then generates the corresponding game description $y$. The query $x$ is a natural language sentence describing the content or rules of a game. Our goal is to make the game description $\hat{y}$, generated by the LLM, as close as possible to the ground-truth game description $y$. In this study, we employ Ludii GDL, a Game Description Language (GDL) commonly used in recent research. In this study, we use Ludii GDL as one of the forms of GDL. This is because Ludii GDL is a versatile framework capable of describing a wide variety of classical games, and it has been widely employed in recent research on GDL [16, 17, 12, 18, 19, 11].
@@ IV-A Training Procedure
Our training procedure consists of two steps. First, we perform SFT using paired data of queries and corresponding game descriptions. Then, we conduct RLFT starting from the SFT-trained model.
SFT in the first step allows llms to avoid unstable outputs commonly encountered in the initial stages of RL. For example, to simplify the extraction of generated programs, we instruct the model to output in a specific format, such as <program>...</program>. Models without SFT often ignore the format instructions, generating extraneous text or incorrect formats like ```xml (program)...```. By resolving these formatting issues through the SFT process, the subsequent RL step can focus solely on improving the quality of the game description.
@@ IV-B Reward Modeling for GDG
Grammar rewards measure how much of the output $\hat{y}$ is grammatically valid according to the GDL grammar. Using an Earley parser [34], we parse $\hat{y}$ from the beginning based on the GDL grammar and obtain the longest substring $\hat{y}_{\mathrm{valid}}$ that is grammatically correct. Let $L_{\hat{y}}$ be the string length of $\hat{y}$ and $L_{\hat{y}_{\mathrm{valid}}}$ be the string length of $\hat{y}_{\mathrm{valid}}$. The grammar reward $r_{g}$ is calculated as follows:
@@
Concept rewards evaluate the similarity between the game features of the predicted output $\hat{y}$ and those of the ground truth $y$. An overview of the calculation process for the concept reward is shown in Fig. 3. First, we determine whether the output has the necessary playable functionality. If rules fail to function properly—such as when a player cannot move their pieces—the output is considered non-functional, and the reward $r_{c}$ is set to 0. If the output is functional, we proceed to the next step. In the next step, we compute the concept values for the output $\hat{y}$ and compare them with the concept values of $y$. Concept values represent features of a game and were introduced in [12]. Although there are hundreds of concept values, inspired by GAVEL, we evaluate the following five items.
@@
    DecisionMoves $c_{1}$: Percentage of moves where there was more than one possible move.
@@
Here, $w$ represents weights, all of which are set to 0.18. If the five concept values cannot be computed—for example, when the game is functional but automatic playout calculation exceeds the timeout limit—then $r_{c}$ is set to 0.1. This corresponds to the case where $c_{1},\dots,c_{5}$ are all equal to 1.0; in other words, when the game features of $\hat{y}$ and $y$ significantly differ. Note that the timeout was set to 180 seconds.
@@ V-A Datasets
The query $x$ consists of metadata provided by the Ludii game system [36], specifically description and rules. The description gives an overview of the game, while the rules detail the game’s specific rules. Following prior work [29, 11], and to enhance the dataset’s generality, we use a game description $y$ in which the game-specific functions defined within each game are fully expanded. After expansion, the game descriptions rely exclusively on primitive functions, omitting Ludii’s meta-language features such as definitions, options, rulesets, ranges, and constants.
@@ V-B Comparison Methods
  • RLGDG (ours): Our proposed approach, which employs an LLM first fine-tuned with SFT and then further fine-tuned with RL). In RLFT, we use the same dataset as the SFT step, as described in Subsection V-A.
@@ VI-A Evaluation Metrics
To evaluate the generated game descriptions, we employ the following metrics based on prior work [11]:
  • Compilability: The proportion of games that can be successfully parsed and compiled by the Ludii game engine. Only games with compilability are further evaluated for functionality. This score is normalized to a range from 0 to 100.
  • Functionality: The proportion of games considered playable, as described in Section IV-B. Specifically, we assess whether the generated output has sufficient functionality to allow proper gameplay. If the game’s rules fail to operate correctly, the game is considered non-functional. This score is normalized to a range from 0 to 100.
@@
  • Normalized Concept Distance (NCD): A metric to measure the similarity of game features between generated outputs and ground truth. Based on [12, 16], games are represented as concept value vectors, and their cosine distance is calculated to determine NCD. The concept value vector includes semantic features and behavior data from random playouts, such as the proportion of board positions used at least once, or the proportion of turns where at least one legal move exists. It also includes the five concept rewards described in Section IV-B. For ground truth games, 50 playouts are executed, while for generated games, 10 playouts are performed. For non-functional games where the concept distance cannot be calculated, NCD is set to 1. The average NCD is computed to evaluate the quality of game description generation.
Experiments are conducted using three different random seeds, and the mean and standard error for each metric are reported.
@@ VIII Discussion and Conclusion
In this study, we propose Reinforcement Learning-based fine-tuning of LLMs for Game Description Generation (RLGDG). Existing approaches have primarily focused on improving grammatical accuracy; however, our method simultaneously enhances both grammatical correctness and conceptual fidelity to game concepts. Specifically, we introduce grammar and concept rewards and adopt a two-stage training strategy that applies RL after Supervised Fine-Tuning (SFT). Experimental results demonstrated that our proposed method achieved superior performance compared to baseline methods in terms of both grammatical accuracy and conceptual fidelity. However, improvements are limited in categories with insufficient training data. Future research directions for training data include synthetic data generation [40], data augmentation using evolutionary algorithms [29], and leveraging larger-scale language models. These studies are expected to generate high-quality game descriptions from natural language, thereby supporting designers and engineers in AI-driven game development.

### v2

@@ Abstract
Game Description Generation (GDG) is the task of generating a game description written in a Game Description Language (GDL) from natural language text. Previous studies have explored generation methods leveraging the contextual understanding capabilities of Large Language Models (LLMs); however, accurately reproducing the game features of the game descriptions remains a challenge. In this paper, we propose reinforcement learning-based fine-tuning of LLMs for GDG (RLGDG). Our training method simultaneously improves grammatical correctness and fidelity to game concepts by introducing both grammar rewards and concept rewards. Furthermore, we adopt a two-stage training strategy where Reinforcement Learning (RL) is applied following Supervised Fine-Tuning (SFT). Experimental results demonstrate that our proposed method significantly outperforms baseline methods using SFT alone. Our code is available at https://github.com/tsunehiko/rlgdg
@@ I Introduction
Recently, there has been increased interest in Game Description Generation (GDG) [9, 10]. GDG focuses on generating game descriptions from natural language texts, making it easier for non-experts to participate in game design. In GDG tasks, In-Context Learning (ICL) with Large Language Models (LLMs) [11] has shown great promise. LLMs excel in understanding textual contexts and can perform tasks with limited domain knowledge based on a few examples provided in prompts. For example, Hu et al. [9] demonstrated that enriching prompts with explanations of GDL notation and examples of game descriptions improves the accuracy of generated outputs. Additionally, Grammar-based Game Description Generation (GGDG) [10] proposed an iterative decoding method guided by GDL grammar rules, significantly enhancing grammatical correctness. However, the iterative improvements in GGDG have restricted grammatical accuracy and do not consider the actual gameplay behavior and features obtained through simulation.
@@ II-A Game Description Language
Game Description Language (GDL) is a specialized language for describing specific games. In 2005, GGP-GDL [1] was introduced for General Game Playing, aiming to develop artificial intelligence agents capable of adapting to various games. Video Game Description Language (VGDL) [2] has been developed to represent rules and levels of 2D sprite-based games, currently modeling as many as 195 different games. Regular Boardgames (RBG) [3] is another language that combines high-level language features with low-level descriptions, enabling the representation of complex board games. The Ludi system [4], incorporating evolutionary game design, successfully led to the development of the commercially successful game “Yavalath.” Moreover, Ludii [5] is a system adopting the “ludemic approach,” allowing the decomposition and description of game components at a conceptual level. This enables Ludii to represent over 1,000 traditional games, including board games, card games, dice games, and tile games. Given the broad representational capability and versatility of Ludii GDL, we primarily use Ludii GDL for our research.
@@ II-B Large Language Models in Games
Since the emergence of ChatGPT [20], large language models (LLMs) have garnered significant attention, prompting exploration into diverse applications within the field of game AI [21, 22]. For example, several methods [23, 24] have been proposed for generating 2D tile-based game levels by fine-tuning GPT-2 [11]. Additionally, research is progressing on prompt-based LLM methods for level generation [25] and quest generation in role-playing games [26]. Dreamcraft [27] further demonstrates a technique to create 3D game objects within Minecraft from textual prompts. Li et al. [28] introduces a virtual pet game that achieves real-time gameplay even with smaller LLMs by employing a domain-specific distillation approach.
On the other hand, research has also advanced in automatically generating game rules and descriptions using LLMs and GDL. GAVEL [29] employs an LLM fine-tuned on Ludii game descriptions as a mutation operator in evolutionary search, aiming to create novel games. The objective of GAVEL differs from our research, which focuses on generating game descriptions that maintain consistency with natural language text. Hu et al. [9] proposes a method for generating both rules and levels in VGDL using LLMs. LLMaker [30] improves content consistency by utilizing function calling, but its creativity is limited to the scope of the defined functions. GGDG [10] utilizes iterative decoding based on the grammar of Ludii GDL to enhance the grammatical correctness of generated game descriptions. While these previous studies primarily focus on novelty or grammatical correctness of game descriptions, our research differs by emphasizing the improvement of game characteristics of generated game descriptions, bringing them closer to the ground truth through reinforcement learning to enhance LLMs.
@@ II-C RL for LLMs
RL-based fine-tuning of LLMs [31, 14, 13, 15] has been attracting attention. Reinforcement Learning with Human Feedback (RLHF) [31] aligns model behaviors with human preferences by using human feedback as rewards. Recent methods [14, 13] have explored accuracy-based reward functions without human feedback, notably improving logical reasoning tasks such as mathematics and programming. RL-finetuned models are known to exhibit enhanced reasoning capabilities [14]. Additionally, advancements in RL algorithms led to the development of Group Relative Policy Optimization (GRPO) [32], a variant of Proximal Policy Optimization (PPO) [33] optimized for fine-tuning LLMs. In this study, we focus on GDG and define a reward function based on both grammatical correctness and conceptual relevance to the game in generated descriptions, employing GRPO to fine-tune the LLM.
Figure 2: An example of GDG for the game Tic-Tac-Toe. $x$ is text that explains games in natural language. $\hat{y}$ is the game description generated by the LLM in Ludii GDL, a Game Description Language.
@@ III Problem Setting
In this section, we define GDG, the primary task addressed in this paper. As shown in Fig. 2, the task involves providing a query $x$ to a large language model (LLM), which then generates the corresponding game description $y$. The query $x$ is a natural language sentence describing the content or rules of a game. Our goal is to make the game description $\hat{y}$, generated by the LLM, as close as possible to the ground-truth game description $y$. The acquisition of query $x$ and ground truth $y$ is described in Section V-A.
@@ IV-A Training Procedure
Our training procedure consists of two steps. First, we perform SFT using the paired data of queries and corresponding game descriptions. Then, we conduct RLFT starting from the SFT-trained model.
SFT in the first step allows LLMs to avoid unstable outputs commonly encountered in the initial stages of RL. For example, to simplify the extraction of generated programs, we instruct the model to output in a specific format, such as <program>...</program>. Models without SFT, especially those with smaller parameter sizes such as 1.5B, often ignore the format instructions, generating extraneous text or incorrect formats like ```xml (program)...```. By resolving these formatting issues through the SFT process, the subsequent RL step can focus solely on improving the quality of the game description.
@@ IV-B Reward Modeling for GDG
Grammar rewards measure how much of the output $\hat{y}$ is grammatically valid according to the GDL grammar. Using the Earley parser implemented in Ludii [34], we parse $\hat{y}$ from the beginning based on the GDL grammar and obtain the longest grammatically correct substring $\hat{y}_{\mathrm{valid}}$. Let $L_{\hat{y}}$ be the string length of $\hat{y}$ and $L_{\hat{y}_{\mathrm{valid}}}$ be the string length of $\hat{y}_{\mathrm{valid}}$. The grammar reward $r_{g}$ is calculated as follows:
@@
Concept rewards evaluate the similarity between the game features of the predicted output $\hat{y}$ and those of the ground truth $y$. An overview of the calculation process for the concept reward is shown in Fig. 3. First, we determine whether the generated game is functional. To clarify this, we introduce two notions: Compatibility indicates whether the Ludii game engine can parse and compile the game without errors, and Functionality indicates whether the compiled game works well enough to be played. If rules fail to function properly, such as when a player cannot move their pieces, the output is considered non-functional. Only compilable games can be evaluated for functionality. A functional game proceeds to the next evaluation step; otherwise, the reward $r_{c}$ is set to 0. In the next step, we compute the concept values for the output $\hat{y}$ and compare them with the concept values of $y$. Concept values represent features of a game and were introduced in [12]. Although there are hundreds of concept values, as an initial exploration of incorporating game concepts into RL, we evaluate the following five items, taking inspiration from GAVEL [29].
@@
    DecisionMoves $c_{1}$: Percentage of terns where there was more than one possible move.
@@
Here, $w$ denotes the weights. When the game is functional but the game features of $\hat{y}$ and $y$ differ greatly—that is, when $c_{1},\dots,c_{5}$ are all equal to $1.0$—we grant a small reward of $r_{c}=0.1$ for simply functioning. We implement this by setting every $w$ to $0.18$. Furthermore, if the five concept values cannot be computed—for example, when the game is functional but the automatic playout calculation exceeds the timeout limit—we also set $r_{c}$ to $0.1$. Note that the timeout is set to $180$ seconds.
@@ V-A Datasets
The query $x$ consists of metadata provided by the Ludii game system [36], specifically description and rules. “Description” gives an overview of the game, while “Rules” detail the game’s specific rules. Following prior work [29, 10], and to enhance the dataset’s generality, we use a game description $y$ in which the game-specific functions defined within each game are fully expanded. After expansion, the game descriptions rely exclusively on primitive functions, omitting Ludii’s meta-language features such as definitions, options, rulesets, ranges, and constants.
@@ V-B Comparison Methods
  • RLGDG (ours): Our proposed approach, which employs an LLM first fine-tuned with SFT and then further fine-tuned with RL. In RLFT, we use the same dataset as the SFT step, as described in Section V-A.
@@ VI-A Evaluation Metrics
To evaluate the generated game descriptions, we employ the following metrics based on prior work [10]. For details on Compilability and Functionality, see Section IV-B.
  • Compilability: The proportion of games that can be successfully parsed and compiled by the Ludii game engine. This score is normalized to a range from 0 to 100.
  • Functionality: The proportion of games considered playable. This score is normalized to a range from 0 to 100.
@@
  • Normalized Concept Distance (NCD): A metric to measure the similarity of game features between generated outputs and ground truth. Based on [12, 16], games are represented as concept value vectors, and their cosine distance is calculated to determine NCD. The concept value vector includes semantic features and behavior data from random playouts, such as the proportion of board positions used at least once, or the proportion of turns where at least one legal move exists. It also includes the five concept rewards described in Section IV-B. For ground truth games, 50 playouts are executed, while for generated games, 10 playouts are performed. For non-functional games where the concept distance cannot be calculated, NCD is set to 1. The average NCD is computed to evaluate the quality of GDG.
Experiments are conducted using three different random seeds following [10], and the mean and standard error for each metric are reported.
@@ VIII Discussion and Conclusion
In this study, we propose Reinforcement Learning-based fine-tuning of LLMs for Game Description Generation (RLGDG). Existing approaches have primarily focused on improving grammatical accuracy; however, our method simultaneously enhances both grammatical correctness and conceptual fidelity to game concepts. Specifically, we introduce grammar and concept rewards and adopt a two-stage training strategy that applies RL after Supervised Fine-Tuning (SFT). Experimental results demonstrated that our proposed method achieved superior performance compared to baseline methods in terms of both grammatical accuracy and conceptual fidelity. However, improvements are limited in categories with insufficient training data. Future research directions for training data include synthetic data generation [40], data augmentation using evolutionary algorithms [29], and leveraging larger-scale language models. Moreover, extending the concept reward from five to the full set of concept values remains a key target. These studies are expected to generate high-quality game descriptions from natural language, thereby supporting designers and engineers in AI-driven game development.
Acknowledgments
This work was supported by JST, ACT-X Grant Number JPMJAX23CE, Japan.